OpenStudy (firejay5):

10^ 7 = 10,000,000 to a logarithmic statement; I said log(10000000) = 7 is this correct?

OpenStudy (perl):

correct.

OpenStudy (firejay5):

thank you perl just checking if I did it right.

OpenStudy (firejay5):

what about log a = b in an exponential statement?

OpenStudy (perl):

$$ \LARGE{ \log _{10} (a) = b \iff 10^b = a } $$

OpenStudy (perl):

if the base is not given, you can assume it is 10 (usually).

OpenStudy (firejay5):

so is: 10^b = a the same thing pretty much

OpenStudy (firejay5):

thank you so much!

OpenStudy (perl):

right :)